# Post-detection inference for sequential changepoint localization

### v3

@@ 3.2 A partial extension to composite pre-change
$\exists F_{0}^{*}\in\mathcal{P}_{0}$ such that $\inf_{F_{0}\in\mathcal{P}_{0}}\mathbb{P}_{F_{0},\infty}(\tau\geq t)\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t)$, $\forall t\in\mathbb{N}$.
This assumption posits the existence of a “least favorable” distribution $F_{0}^{*}$, in the sense that the detection algorithm is most prone to false alarms under $F_{0}^{*}$. We show that it holds for monotone likelihood ratio families with CUSUM/SR type change detectors when the pre- and post-change parameter spaces, $\Theta_{0}$ and $\Theta_{1}$, are ordered, e.g., $\Theta_{0}=(-\infty,a],\Theta_{1}=(b,\infty)$, for some $a<b$.  [^5] For detection algorithms with bounded PFA, one can avoid Assumption 3.3 and obtain confidence sets with coverage guarantees without any additional assumption (i.e., for arbitrary $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$); see Section F of the Supplementary.
@@
Suppose that $\{F_{\theta}:\theta\in\Theta\}$ is some monotone likelihood ratio family with respect to some sufficient statistic for the parameter $\theta$ and $\sup\Theta_{0}\leq\inf\Theta_{1}$. For the CUSUM type stopping time $\tau$ defined in (C.12), and for any $\theta\leq\theta^{\prime}\leq\sup\Theta_{0}$, $\mathbb{P}_{\theta,\infty}(\tau\geq t)\geq\mathbb{P}_{\theta^{\prime},\infty}(\tau\geq t),$ and hence, $\theta_{0}^{*}=\sup\Theta_{0}$ satisfies Assumption 3.3. Similarly, if $\inf\Theta_{0}\geq\sup\Theta_{1}$, then $\theta_{0}^{*}=\inf\Theta_{0}$ satisfies Assumption 3.3.
For each $j=1,\dots,N$, we now simulate a sequence of i.i.d. data points $Z_{1}^{(j)},Z_{2}^{(j)},\dots\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{0}^{*}$ until either the detection algorithm triggers or the sample size reaches $\tau$—whichever occurs first. Let $\tau_{j}^{*}:=\min\{\tau,\mathcal{A}(\{Z_{n}^{(j)}\}_{n})\}$ denote the stopping time for $\{Z_{n}^{(j)}\}_{n}$. We estimate the probability $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$ using
$r_{t}^{*}=\frac{1}{N}\sum_{j=1}^{N}\mathds{1}(\tau_{j}^{*}\geq t).$ (3.7)
Define the confidence set as
@@
Suppose that the pre-change class $\mathcal{P}_{0}$ and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfy Assumption 3.3 and $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0,\forall F_{0}\in\mathcal{P}_{0}$. Then, for any $\alpha\in(0,1),$ $T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, the confidence set $\mathcal{C}$ in (3.8) satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
@@ 4 A simulation-based approach for parametric pre- and post-change settings
In this section, we introduce an alternative approach that, while restricted to the parametric setting, allows observations to be dependent, and often yields tighter confidence sets than the method proposed in the previous section, as shown by our simulations later. We begin with the most basic case, where the pre- and post-change distributions are known, and then extend it to more general parametric settings subsequently.
@@ 4.1 Known pre- and post-change distributions
We develop a simulation-based scheme to approximate the exact $\alpha$-th quantile of the distribution of the test statistic $M_{t}$. Here, $M_{t}$ need not be of the form (3.4), it can be any test statistic based on $X_{1},\cdots,X_{\tau}$. Assuming known pre- and post-change distributions, we draw $B$ many independent streams of data under $H_{0,t}$. For the $j$-th sequence, denoted as $\{X_{n}^{j}\}_{n}$, we simulate data points until a change is detected at $\tau_{j}^{\prime}=\mathcal{A}(\{X_{n}^{j}\}_{n})$ or until a fixed time $L$, whichever happens first. Then, we compute the test statistic $M_{t,L}^{j}$ as described in Section 4.1.1, based on $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$. Our confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$:
@@
where $r_{t}$ is as defined in (3.7). Algorithm 1 contains an overview of the method.
@@
For any fixed $\alpha\in(0,1)$, $B,N,T\in\mathbb{N}$, ${}L\in\mathbb{N}\cup\{\infty\}$, and change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}_{1}$ defined in (4.1) satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-{\alpha}.$
@@
4.1.1 Subtleties regarding Subroutine $\mathcal{L}$
The problem with $L=\infty$ is that the stopping time may never be reached for some sampled sequence. One can set $L=\infty$ if $\mathcal{A}$ is known to satisfy $\mathbb{P}_{T}(\tau<\infty)=1$, for all $T\in\mathbb{N}$ (if there is a change, $\mathcal{A}$ stops almost surely). This is a very mild assumption because any reasonable change detector should eventually raise an alarm if there is a change; indeed, for example, this holds for all algorithms with finite ARL. If $\mathcal{A}$ is not known to satisfy this condition, one can simply pick $L<\infty$. If $L<\infty$ and the $j$th sequence stopped because $\mathcal{A}$ raised an alarm at or before $L$, then we define $M^{j}_{t,L}$ as usual. If $\mathcal{A}$ had not stopped by time $L$, we then define $M^{j}_{t,L}=\infty$ to ensure a lower bound on the original test statistic (if we could compute it). For $L<\infty$,
$M^{j}_{t,L}:=\begin{cases}-\infty,&\text{if }t>\tau_{j}^{\prime},\\ M_{t}^{(\tau_{j}^{\prime})}(X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}}^{j}),&\text{if }t\leq\tau_{j}^{\prime}\leq L,\\ \infty,&\text{if }t\leq\tau_{j}^{\prime}\text{ and }\tau_{j}^{\prime}>L.\\ \end{cases}$ (4.2)
Remark 4.3.
The value of $L$ does not need to be a fixed constant; it can be set adaptively depending on the observed data sequence, for instance, one can choose $L$ to be $2\tau$.
@@ 4.2 Adaptive confidence set for known pre-change distribution and parametric post-change class
[…]
@@
We construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, which have $1-\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$ coverage, where $r_{t}$ is as defined in (3.5). Despite $r_{t}$ being random, it turns out that marginalizing over $r_{t}$ yields that the CS has coverage $\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$; we formalize this more explicitly in the proof of the Theorem 4.4.
Fix $B\in\mathbb{N}$. For each $j=1,\cdots,B$, suppose we have a sequence $\{X^{j}_{n}(\theta_{0},t,\theta)\}_{n}$ $\sim\mathbb{P}_{\theta_{0},t,\theta}$. Note that for a fixed $t$ and $\theta$, the sequences must be independent across $j$. However, for any fixed $j$, the i.i.d. sequences $\{X^{j}_{n}(\theta_{0},\theta)\}_{n}$ are allowed to be dependent across different values of $t$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. A specific method for generating these sequences is detailed in Section B of the Supplementary. We draw the data until a change is detected at $\tau_{j,t}^{\theta}$ using $\mathcal{A}$ or until a fixed time $L$, whichever occurs first. Then, we compute the same test statistics truncated at $L$, based on $\tau_{j,t}^{\theta}$ and $X^{j}_{1}(\theta_{0},t,\theta),\cdots,X^{j}_{\tau_{j,t}^{\theta}\wedge L}(\theta_{0},t,\theta)$, which is denoted as $M^{j}_{t,L}(\theta)$. For $L<\infty$,
@@
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\}$ (4.4)
Algorithm 2 summarizes the method. Note that, even when $\mathcal{S}^{\prime}_{\tau-t+1}$ is an infinite set, the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ can be approximated by discretizing $\mathcal{S}^{\prime}_{\tau-t+1}$ into a grid and evaluating the quantile for values of $\theta$ on this grid, taking their maximum. While this approach is conceptually straightforward, it is computationally intensive; related experiments are in Section E of the Supplementary. Moreover, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient and does not require any approximation. Details are provided in Section B of the Supplementary, along with two concrete examples of the Gaussian and Poisson mean change problems.
@@
Let $\mathcal{C}_{1}$ be the confidence set for changepoint $T\in\mathbb{N}$ as defined in (4.4) for some $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N}$, ${}L\in\mathbb{N}\cup\{\infty\}$. Then for any given detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$ and any $\theta_{1}\in\Theta_{1}$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta.$

### v4

@@ 3.2 An extension to composite pre-change
$\exists P_{0}^{*}\in\mathcal{P}_{0}$ such that $\inf_{P_{0}\in\mathcal{P}_{0}}\mathbb{P}_{P_{0},\infty}(\tau\geq t)\geq\mathbb{P}_{P_{0}^{*},\infty}(\tau\geq t)$, $\forall t\in\mathbb{N}$.
We justify this assumption by appealing to Huber’s notion of “least favorable distribution pair” [13] in terms of risk, where the risk $R$ for a hypothesis test $\phi$ between $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$ is defined as
$R(P_{0},\phi)=C_{0}P_{0}(\phi\text{ accepts }\mathcal{P}_{1}),\penalty 10000\ \penalty 10000\ R(P_{1},\phi)=C_{1}P_{1}(\phi\text{ accepts }\mathcal{P}_{0}).$ (3.8)
The formal definition of a least favorable distribution pair is given below.
Definition 3.6.
A pair $(P_{0}^{*},P_{1}^{*})\in\mathcal{P}_{0}\times\mathcal{P}_{1}$ is called least favorable distribution (LFD) pair in terms of risk $R$ for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$, if for every likelihood ratio test $\phi$ between $P_{0}^{*}$ and $P_{1}^{*}$, $R(P_{j}^{*},\phi)\geq R(P_{j},\phi),\penalty 10000\ \text{ for all }P_{j}\in\mathcal{P}_{j},\penalty 10000\ j=0,1.$
The following proposition establishes that the existence of an LFD pair is sufficient for our assumption to hold.
@@
$(P_{0}^{*},P_{1}^{*})$ is a least favourable distribution pair in terms of risk for testing $\mathcal{P}_{0}$ vs. $\mathcal{P}_{1}$, then Assumption 3.5 holds with $F^{*}=P^{*}$, and $\tau$ being the stopping time of a CUSUM/SR change detector with the likelihood ratio of $P_{1}^{*}$ to $P_{0}^{*}$.
LFD pairs are known to exist for various parametric and nonparametric models, e.g., the monotone likelihood ratio (MLR) families, and Huber’s robust models (including $\epsilon$ neighborhood model, total variation model, and more generally with 2-alternating capacities [14]).
Suppose we have estimators $r_{t}^{*}$ of $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$, for each $t$ (which can be constructed using simulations from $F^{*}_{0}$ as described in Section 3.1). Now, we analogously define the confidence set as
@@
Suppose that the pre-change class $\mathcal{P}_{0}$ and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfy Assumption 3.5 and $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0,\forall F_{0}\in\mathcal{P}_{0}$. Then, for any $\alpha\in(0,1),$ $T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, the confidence set $\mathcal{C}$ defined in (3.9), constructed using any data-independent unbiased estimator $r_{t}^{*}$ of $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$
We now analyze the size of the confidence set, generalizing Theorem 3.3 to the case where an LFD pair $(F_{0}^{*},F_{1}^{*})$ exists for $\mathcal{P}_{0}$ vs $\mathcal{P}_{1}$. We consider the confidence set in (3.9) with $r_{t}^{*}=\frac{1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}^{*}\geq t)}{N+1}$, $M_{t}$ as defined in (3.4) with $\hat{T}$ as defined in (3.3) using $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})},$ and $S_{n}^{(t)}=\prod_{i=n}^{t-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}$. We redefine $p_{t},\rho_{0}$ and $\rho_{1}$ in terms of the LFD pair: $p_{t}=\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t)$, $\rho_{0}(s)=\mathbb{E}_{F_{0}^{*}}\left(\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\right)^{s}$ and $s_{0}=\arg\min_{s\in(0,1)}\rho_{0}(s)$ and $\rho_{1}(s)=\mathbb{E}_{F_{1}}\left(\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\right)^{s}$ and $s_{1}=\arg\min_{s\in(0,1)}\rho_{1}(s)$. Finally, we call the stopping time $\tau$ well-behaved if it satisfies
$\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ $\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq t),\forall t\in\mathbb{N},\forall F_{0}\in\mathcal{P}_{0};$ (3.10a)
$\mathbb{P}_{F^{*}_{0},T,F^{*}_{1}}[\tau\leq t]$ $\geq\mathbb{P}_{F^{*}_{0},\infty}[\tau\leq t],\forall t,T\in\mathbb{N}\penalty 10000\ \text{ and }$ (3.10b)
$\mathbb{P}_{F_{0},T,F_{1}}[\tau\leq t]$ $\geq\mathbb{P}_{F^{*}_{0},T,F^{*}_{1}}[\tau\leq t],\forall t,T\in\mathbb{N},\text{ and }\forall(F_{0},F_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}.$ (3.10c)
Note that (3.10b) is analogous to (3.6), while (3.10c) requires that the stopping time $\tau$ under change from $F_{0}^{*}$ to $F_{1}^{*}$ is stochastically larger than $\tau$ under change from $F_{0}$ to $F_{1}$, for any $(F_{0},F_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}.$ To justify that (3.10c) is sensible, we note that the LFD pair $(F_{0}^{*},F_{1}^{*})$ is the closest pair within $\mathcal{P}_{0}\times\mathcal{P}_{1}$ (according to various metrics or divergences such as Kullback-Leibler), i.e., it renders the change detection problem most difficult. Therefore, it is reasonable to assume that $\tau$ under change from $F_{0}^{*}$ to $F_{1}^{*}$ is stochastically largest. (3.10a) can be justified similarly because $F_{0}^{*}$ is the “closest” element of the pre-change class to the post-change class.
Theorem 3.9.
In the setting described above, for any change detection method with stop time $\tau$ that is well-behaved in the sense of (3.10a) and $p_{T}\neq 0$,
$\limsup_{N\to\infty}\sup_{\begin{subarray}{c}F_{0}\in\mathcal{P}_{0}\\ F_{1}\in\mathcal{P}_{1}\end{subarray}}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\frac{1-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}},$ (3.11)
where $\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}}:=\displaystyle\sup_{\begin{subarray}{c}F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}\end{subarray}}\Delta_{F_{0},T,F_{1}}$. Additionally, if the stopping time $\tau$ is well-behaved in the sense of (3.10b) and (3.10c) as well, the last term in (3.11) can be improved to $\min\left\{\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right),\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}}\right\}.$
@@ 4 A simulation-based approach for parametric pre- and post-change settings
In this section, we introduce an alternative approach that, while restricted to the parametric setting, allows observations to be dependent, and often yields tighter confidence sets than the method proposed in the previous section, as shown by our simulations later. We begin with the most basic case, where the pre- and post-change distributions are known, and then extend it to more general parametric settings.
@@ 4.1 Known pre- and post-change distributions
We develop a simulation-based scheme to approximate the exact $\alpha$-th quantile of the distribution of the test statistic $M_{t}$. Here, $M_{t}$ need not be of the form (3.4), it can be any test statistic based on $X_{1},\cdots,X_{\tau}$. Assuming known pre- and post-change distributions, we draw $B$ many independent streams of data under $H_{0,t}$. For the $j$-th sequence, denoted as $\{X_{n}^{j}\}_{n}$, we simulate data points until a change is detected at $\tau_{j}^{\prime}=\mathcal{A}(\{X_{n}^{j}\}_{n})$ or until a fixed time $L$, whichever happens first. Then, we compute a test statistic $M_{t,L}^{j}$, based on $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$ (details are described in Supplementary Section B). Our confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$:
@@
where $r_{t}$ is any unbiased estimator of $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$; see Algorithm 1.
@@
For any fixed $\alpha\in(0,1)$, $B,N,T\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$, and change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}_{1}$ defined in (4.1), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-{\alpha}.$
@@ 4.2 Adaptive confidence set for known pre-change distribution and parametric post-change class
Since the post-change parameter is unknown, for each $t$, we estimate it based on $X_{t},\cdots,X_{\tau}$. However, the sample size used for the estimate might be too small to have a reliable estimate. To address this issue, we would use interval estimates instead of point estimates, allowing us to achieve a predetermined control over the error. Notably, conventional confidence intervals are invalid here, since $\tau$ is random. Therefore, we rely on confidence sequences [4, 12]. A confidence sequence having coverage 1c1-c is defined as a sequence of confidence sets $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n\in\mathbb{N}}$, satisfying the uniform guarantee:
@@
We construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, which have $1-\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$ coverage, where $r_{t}$ is any unbiased estimator of $\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$.
Fix $B\in\mathbb{N}$. For each $j=1,\cdots,B$, suppose we have a sequence $\{X^{j}_{n}(\theta_{0},t,\theta)\}_{n}$ $\sim\mathbb{P}_{\theta_{0},t,\theta}$. Note that for a fixed $t$ and $\theta$, the sequences must be independent across $j$. However, for any fixed $j$, the i.i.d. sequences $\{X^{j}_{n}(\theta_{0},\theta)\}_{n}$ are allowed to be dependent across different values of $t$ and $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. A specific method for generating these sequences is detailed in Section C of the Supplementary. We draw the data until a change is detected at $\tau_{j,t}^{\theta}$ using $\mathcal{A}$ or until a fixed time $L$, whichever occurs first. Then, we compute the same test statistics truncated at $L$, based on $\tau_{j,t}^{\theta}$ and $X^{j}_{1}(\theta_{0},t,\theta),\cdots,X^{j}_{\tau_{j,t}^{\theta}\wedge L}(\theta_{0},t,\theta)$, which is denoted as $M^{j}_{t,L}(\theta)$. For $L<\infty$,
@@
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\}.$ (4.3)
Algorithm 2 summarizes the method. Note that, even when $\mathcal{S}^{\prime}_{\tau-t+1}$ is an infinite set, the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ can be approximated by discretizing $\mathcal{S}^{\prime}_{\tau-t+1}$ into a grid and evaluating the quantile for values of $\theta$ on this grid, taking their maximum. While this approach is conceptually straightforward, it is computationally intensive; related experiments are in Section F of the Supplementary. Moreover, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient and does not require any approximation. Details are provided in Section C of the Supplementary, along with two concrete examples of the Gaussian and Poisson mean change problems.
@@
Let $\mathcal{C}_{1}$ be the confidence set for changepoint $T\in\mathbb{N}$ as defined in (4.3), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{\theta_{0},\infty}[\tau\geq t]$, for some $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$. Then for any given detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$ and any $\theta_{1}\in\Theta_{1}$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta.$